OpenStudy (anonymous):

What is the equation in point-slope form of the line passing trhough (0,2) and (1,5)? y - 5 = 3(x-1) y - 2 = -3(x-0) y +2 = 3(x+1) y + 5 = -3(x+1) I solved the equation using point slope form for a linear equation and I did it twice, one for each set of coordinates and I got A and B. I just don't know which one to choose that'll be right since I got both answers. Help please?

OpenStudy (texaschic101):

find the slope first.....slope(m) = (y2 - y1) / (x2 - x1) (0,2).....x1 = 0 and y1 = 2 (1,5)...x2 = 1 and y2 = 5 slope(m) = (5 - 2) / (1 - 0) slope(m) = 3/1 or just 3 Now use point slope form y - y1 = m(x - x1) slope (m) = 3 you can use either of your points using (0,2).....x1 = 0 and y1 = 2 y - y1 = m(x - x1) y - 2 = 3(x - 0) using (1,5)...x1 = 1 and y1 = 5 y - y1 = m(x - x1) y - 5 = 3(x - 1) so your only possible answers are : y - 2 = 3(x - 0) and y - 5 = 3(x - 1) so B is not correct because the 3 is not negative understand ?

OpenStudy (anonymous):

Ops, sorry. I knew that it wasn't B - I meant to say it's either A or C. The answers I got were... y - 2 = 3(x-0) ---> that was for the corrdinates (0,2) y - 5 = 3(x-1) ---> that was for coordinates (1,5) I just don't know what to do know since I can't have two answers.

OpenStudy (texaschic101):

anything with a -3 as the slope will not work. Thus leaving only 2 answer choices...and 1 of them is y + 2....not an answer choice. So A is the only correct one.

OpenStudy (texaschic101):

I mean, in your answers that you arrived at....there is no y + 2...so C is wrong

OpenStudy (anonymous):

Wait a minute. It's A, sorry. I've been looking at numbers for too long. I got mixed up. Thanks for helping me out :)